OpenStudy (anonymous):

NEEEDD HEEELLP! Pure acid is to be added to a 5% acid solution to obtain 95L of 62% solution. What amounts should be used of each? How many liters of 100% pure acid should there be used to make a solution

OpenStudy (anonymous):

@undeadknight26

OpenStudy (aum):

Assume x liters of 100% pure acid is added to y liters of 5% acid solution to get 95 liters of 62% solution. Then, x + y = 95 ---- (1) x + 0.05y = 95*0.62 or x + 0.05y = 58.9 ---- (2) Eliminate y and solve for x from equations (1) and (2).